Set up an equation that can be used to solve the problem. Solve the equation and answer the question asked:
One number is 5 less than 3 times a second number. Their sum is 43. Find the numbers.

Found 2 solutions by jorel1380, stanbon:
Answer by jorel1380(3719)

p+q=43
q=3p-5
p+3p-5=43
4p=48
p=12
q=31..

Answer by stanbon(75887)

One number is 5 less than 3 times a second number.
Their sum is 43. Find the numbers.
-------------------
x = 3y -5
x + y = 43
----
Substitute for "x" and solve for "y":
3y-5 + y = 43
4y = 38
y = 19/2 = 9.5
---------------------
Solve for "x":
x = 3(9.5) - 5
---
x = 28.5 - 5
---
x = 23.5
============
Cheers,
Stan H.
=============